This question is typical on some driver’s license exams: A car moving at 40 km/h skids
15 m with locked brakes.
How far will the car skid with locked brakes
at 100 km/h? Assume that energy loss is due
only to sliding friction.
Answer in units of m.
Don't Round answer.
1
Expert's answer
2019-11-01T11:40:56-0400

Let frictional retardation be "a"

"40\ km\ h^{-1}=11.11\ m\ s^{-1}"

Then,

"11.11^2=2\times a\times 15\\or\\a=4.115\ m\ s^{-2}"

Let s be the skidding distance for 2nd case

Then,

"27.77^2=2\times 53.33\times s\\or\\s=93.75\ m"